The product of two consecutive even integers is 840. Find these numbers
Answer by ewatrrr(24785)

 
Hi,
x(x+2) = 840
x^2 + 2x - 840 = 0
(x+30)(x-28)= 0
(x+30)= 0, x = -30 the two consecutive even integers: -30, -28
0r (x-28)= 0, x = 28 the two consecutive even integers: 28, 30